OpenStudy (anonymous):

1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? (2 points) 2. Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? (2 points) 3.Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points)

OpenStudy (anonymous):

1) Belinda is 32 while Tony is 24.

OpenStudy (triciaal):

@All_Hail_Moses directly from the Code of conduct: Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem.

OpenStudy (anonymous):

can some one please actually help me?

OpenStudy (triciaal):

I will work with you.

OpenStudy (triciaal):

1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. let Tony's age = x and Belinda's age = b can you write 2 equations with x + b ?

OpenStudy (anonymous):

umm b=56-2x+8? then x=56-b right?

OpenStudy (triciaal):

yes x + b = 56 b = 2x + 8

OpenStudy (triciaal):

so you can substitute for b and solve for x

OpenStudy (anonymous):

idk what I did wrong, I got 40 for b

OpenStudy (triciaal):

minor correction umm b=56-2x+8? should be b = 2x + 8

OpenStudy (anonymous):

yeah I know, that is what I used

OpenStudy (triciaal):

I got the same check it's fine to me x = 16 b = 2*16 = 8 = 32 + 8 = 40 check 2 16 + 40 = 56

OpenStudy (triciaal):

typo 2*16 + 8 = 32 + 8 = 40

OpenStudy (triciaal):

ok with part 1?

OpenStudy (triciaal):

Tony is 16 and Belinda is 40

OpenStudy (anonymous):

Oh okay so I was right

OpenStudy (triciaal):

try #2 with the same approach. Use the information given to set up equations then solve simultaneously

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

let S = student and A = adult then S + A = 150 and 28.50S + 45.00A = 4770 I would multiply the 2nd eqn by 10 to get rid of fraction and also factor to work with smaller numbers

OpenStudy (triciaal):

using S = 150 -A in the 2nd eqn 57(150 - A) + 90 A = 9540 solve for A = 30 S = 120 good practice to verify solution 28.50(120) + 45(30) = 3420 + 1350 = 4770 as given

OpenStudy (triciaal):

120 students and 30 adults went to the park

OpenStudy (anonymous):

\(\bbox[5pt, #FFFD00 ,border:5px solid #7500FF]{\it\Huge \huge\color{purple}{{-}} \cal\color{#000000}A\color{#000000}l\color{#000000}l~\color{#000000}H\color{#000000}a\color{#000000}i\color{#000000}l~\color{#000000}M\color{#000000}o\color{#000000}s\color{#000000}e\color{#000000}s}\)